# Reference Coverage Analysis of OpenAlex compared to Web of Science and Scopus

### v2

@@ Abstract
OpenAlex is a promising open source of scholarly metadata, and competitor to the established proprietary sources, the Web of Science and Scopus. As OpenAlex provides its data freely and openly, it permits researchers to perform bibliometric studies that can be reproduced in the community without licensing barriers. However, as OpenAlex is a rapidly evolving source and the data contained within is expanding and also quickly changing, the question naturally arises as to the trustworthiness of its data. In this empirical paper, we will study the reference and metadata coverage within each database and compare them with each other to help address this open question in bibliometrics. In our large-scale study, we demonstrate that, when restricted to a cleaned dataset of 16,788,282 recent publications shared by all three databases, OpenAlex has average source reference numbers and internal coverage, respectively, comparable to both Web of Science and Scopus. We also demonstrate that the comparison of other core metadata covered by OpenAlex shows mixed results, with OpenAlex capturing more ORCID identifiers, fewer abstracts and a similar number of Open Access information per article when compared to both Web of Science and Scopus.
@@
Bibliometrics, Open Scholarly Metadata, Citation Analysis, Open Abstracts, ORCID
@@ 1 Introduction
OpenAlex is a promising alternative to proprietary bibliometric data sources as its permissible licensing creates the potential to support a transformation of research practice towards reproducible bibliometrics. This is being realised in open research policies in academia, for example in December 2023, Sorbonne University has switched from using the Web of Science (WoS) and Clarivate bibliometric tools to OpenAlex and open-source tools [^1] https://www.sorbonne-universite.fr/en/news/sorbonne-university-unsubscribes-web-science. Reproducible bibliometric research is hardly possible with proprietary bibliometric data sources as their licensing terms rule out dissemination of data.
@@
As a widely used open source repository of scholarly metadata, OpenAlex has previously been the subject of research, such as the review by [17], which comparatively analysed various APIs to bibliometric corpora, including API interoperability, characteristics and their use in research practice, and [1] who released a working paper on the migration of scholars which included a comparative study between Scopus and OpenAlex, limited to the coverage of scholars in Western and non-Western countries. However, at this early stage of its development, OpenAlex is a highly dynamic data source whose characteristics are not well studied.
In this research report, we present the results of some investigations into the metadata coverage of OpenAlex, following previous quantitative comparisons of citation coverage of the MAG such as [10], intersections of OpenAlex, WoS and Scopus, such as the study of the MAG against Scopus and other databases by Visser [20], journal coverage analysis of WoS, Scopus and Dimensions, such as the study of [16], and data completeness such as Delgado-Quirós and Ortega’s smaller scale comparison of OpenAlex to other databases [2] and Färber’s tool for comparing author records between databases [4], to provide insights into the suitability of OpenAlex for bibliometrics in its current state.
@@
For this purpose, OpenAlex is compared in this paper with two major proprietary bibliometric data sources, WoS and Scopus. The motivation of this study is to analyse the extent to which OpenAlex can serve as an adequate (or even better) and free alternative to established, proprietary databases for bibliometric research and reporting. In this study, we concentrate on the analysis of the reference coverage and the coverage of some additional metadata fields, specifically abstracts, Open Researcher and Contributor IDs (ORCIDs), and Open Access status of items in all three data sources.
We are aware that these initial assessments are likely to change with further developments, so this report should be understood as reflecting the state as of late 2023.
@@ 1.1 Reference Coverage
References are of central importance for bibliometric databases, as matching them to their target items forms the basis for the calculation of citation metrics. As a first step, we compare average reference counts between the three databases, whereby the basis of the comparison are the complete databases, then subsets of publications with the document type ‘article’, and a shared sub-corpus of publications covered by all three databases. Furthermore, references can also be used for an assessment of the coverage, i.e. the proportion of relevant research publications that are included in the database and accessible to users for analysis [16]. An insufficient or biased coverage of the relevant literature should rule out the use of a database for a particular study.
@@
There are different ways to determine the coverage of a database, for example, the comparison with external lists of relevant sources or publication lists of a sample of representative researchers of the studied fields. One relatively simple way to study literature coverage is calculating the internal reference coverage of a database as a whole or in relation to grouping characteristics, such as disciplines, the literature of particular countries or language communities.
@@
The internal coverage is the proportion of those cited references of a publication set which are themselves covered as source items in the database, out of all cited references in the set. We refer to these as source references and in contrast, to references to items that are not themselves indexed in the database as non-source references (or references to non-source items). As an example, suppose a publication set of interest has 5000 references, including multiple references to the same works, of which 2000 are covered by the same database, then the internal coverage of this literature would be 40%.
@@
A more comprehensive introduction to this concept and an analysis of the internal coverage of disciplines–as an estimator for disciplinary coverage–is available in [11, Chapter 7] and [19]. The great advantage of this type of analysis is that one does not need any external data which may be difficult and costly to collect. This reliance on only the assessed data source itself is also the major disadvantage, as one is limited to the reference data as present in the assessed data source with all its contingencies. Therefore one cannot simply extrapolate from the coverage of cited literature to the coverage of literature segments that were never cited in the source data, possibly as a direct consequence of the source database’s selection criteria. These considerations show why internal reference coverage provides merely a partial and possibly source-biased measurement of coverage.
@@
Nevertheless, when comparing citation index databases, the differences in internal reference collection can be a useful guide when choosing the most suitable data source for a planned bibliometric study. There are no established guidelines for numerical values of coverage proportions required to allow reliable studies to be carried out. But, for example, [11] analysed the combined ISI Citation Indexes (the predecessor of today’s Web of Science) and found that the coverage rate, which is the proportion of references from the 2002 source year that refer to ISI source journals, was highest for Molecular Biology and Biochemistry, at around 90%, followed by human-focused Biological Sciences, Chemistry, Clinical Medicine and Physics and Astronomy. It was vastly lower in the Arts and Humanities and intermediary in the Social Sciences, Mathematics and Engineering.
@@
When using this indicator to compare OpenAlex, WoS and Scopus, we are less interested in an evaluation in absolute values, but rather in assessing how OpenAlex performs in comparison to the two established bibliometric databases. Internal reference coverage depends on the size and possibly the disciplinary profile of a database as well as the accuracy of its reference matching procedure. As OpenAlex is actually much larger than Scopus and WoS (see Table 1) it could be expected that its internal reference coverage is at least not lower than in the latter databases.
@@ 1.2 Open Metadata
The increasing discussion surrounding the open availability of various types of scholarly metadata in bibliometrics is not limited to reference coverage, but expands to other metadata [18]. For instance, the Initiative for Open Abstracts (I4OA [^2] https://i4oa.org/) advocates open abstracts of scholarly works and calls on scholarly publishers to submit them to Crossref, a Digital Object Identifier (DOI) registration agency. However, coverage analyses of Crossref suggest that not all publishers provide open scholarly metadata to Crossref [12, 8]. Another example of essential metadata is the use of ORCIDs to persistently identify authors, helping bibliometricians not only to disambiguate author names, but also to interlink different data from different sources based on the ORCID [6]. As open data sources are essential for OpenAlex, we will expand our analysis to compare abstracts and ORCID coverage at the journal level. Moreover, we will assess the coverage of open access status information between OpenAlex and the proprietary databases WoS and Scopus. In contrast to abstract and author information, all three databases use the same source, the open access discovery service Unpaywall, to retrieve open access status information [3].
2 Data and Methods
2.1 Data
In this section, we describe the data used in this study, and the reasoning for our choices of restrictions and subsets of this data. To enable a fair comparison between OpenAlex, and WoS and Scopus, we have created a ‘Shared Corpus’ containing records common to all three datasets based on an exact DOI match, which have been published between 2015 and 2022, where the DOI is unique to the record in all three databases, i.e. there are no multiple records with the same DOI. In the course of database ingestion, it is ensured that publications only ever have one DOI. In a further step, the references of the publications in the Shared Corpus are restricted to those published between 1996 to 2022.
@@ 2.1 Data
Table 1: Sizes of databases and of the Shared Corpus dataset, with the number of references in each
The versions of the WoS, Scopus and OpenAlex databases used in this study are as follows. The WoS and Scopus data are snapshots from five indexes of the WoS Core Collection (Science Citation Index Expanded, Social Sciences Citation Index, Arts & Humanities Citation Index, Conference Proceedings Citation Index – Science and Conference Proceedings Citation Index – Social Sciences) and the Scopus database, both captured in April 2023. The OpenAlex database is the version released in August 2023, due in both cases to the versioning policy of our data host at FIZ Karlsruhe.
@@
As the Scopus database mainly contains items from 1996 onwards (although since 2015, pre-1996 cited references and backfiles of major publishers have been added [^3] https://blog.scopus.com/posts/breaking-the-1996-barrier-scopus-adds-nearly-4-million-pre-1996-articles-and-more-than-83), and WoS and OpenAlex have had no such restriction, to avoid bias in the computation of source reference counts and internal coverage we further restrict references to those items published between 1996 and 2022.
We include in Table 1 a section on articles published 2015-2022 in all three databases, to illustrate the influence of the time restriction to the size of the Shared Corpus, and to give context to the DOI matching and deduplication work described in Section 2.2.
In the Scopus and WoS databases, pre-computed total ’reference counts’, are delivered by the data providers Elsevier and Clarivate, whereas ’source reference counts’ are computed for each record by our data provider FIZ Karlsruhe. Both databases are expected to contain all references of a given publication, regardless of whether they refer to items contained within or not contained within their databases, i.e. whether they are source and non-source references respectively, and without a fixed time restriction. In contrast, in OpenAlex there only exist source references (see the OpenAlex documentation [14] [^4] https://docs.openalex.org/api-entities/works/work-object#referenced_works–apart from a smaller segment of references to supposedly deleted items) and a ‘source reference count’ has been calculated by FIZ Karlsruhe in our database. This fact explains the empty values for the average total reference counts in Table 3. We therefore have to relate this number to the source reference counts taken from WoS and Scopus.
@@
Figure 1: Venn diagram of the intersection sizes of unique DOIs based in each database on exact DOI match, for records published between 2015 and 2022
@@ 2.2 DOI Match and Deduplication
When constructing the Shared Corpus as described in Section 2.1, we excluded records without a DOI and records where more than one publication item is attributed to the same DOI – as we are virtually not able to decide which item is the correct one for a given DOI in the latter case. These duplicate records account for the removal of 39,481 publications (counted as distinct DOI) in addition to those resulting from the restriction to 2015-2022. This accounts for the difference between the size of the Shared Corpus and the nominal intersection of the three databases between 2015 and 2022.
@@ 2.2.1 Error Margins of the DOI Match
Records with a shared DOI or those without a DOI were excluded from the DOI matching step in the construction of the Shared Corpus. In Table 2, which focuses on all publications in the three databases which are published between 2015 and 2022, it can be seen that Scopus has a significantly larger number of DOIs with multiple records associated with it. Altogether, OpenAlex has the greatest number of records without DOI, then WoS and Scopus.
As records without a DOI are not matched in our analyses there is a significant underestimation of the total size of the databases as portrayed in Figure 1, similarly records which have a shared DOIs are counted once. The error margin for the number of records compared to DOIs can be inferred from the data in Table 2.
@@
2.3 Methodology
To evaluate the reference and source reference coverage of WoS and Scopus against OpenAlex, we first used the reported reference counts and pre-calculated source reference counts as described in Section 2.1. The average total reference count and source reference count was computed for: each database, for records marked as ‘article’ (or comprising the document type ‘article’ alongside other type markings, in the case of Scopus and WoS) and for the Shared Corpus resulting from the DOI match (publication years 2015-2022). These numbers were replicated by counting the total number of references and records in each database and taking the ratio (‘references per record’), as reported in Table 4. Then for the final results, queries were created to calculate and average the number of references with reference publication year 1996 to 2022, and the number of references that are linked to source items and publication years 1996 to 2022.
To determine metadata coverage (as detailed in Section 3.3), we also used the Shared Corpus as described in Section 2.1. Here, we restrict to publication items published in journals. For this purpose, the publication type categorisations of Web of Science and Scopus were used and the OpenAlex publications were assigned to these via the DOI comparison of the Shared Corpus, so that OpenAlex could be compared bilaterally with the other two databases. We then specifically compared the coverage of abstracts, ORCIDs and Open Access (OA) status information by assessing whether items have (at least one) of these and aggregated by journal, that is, for each journal, a publication record was counted if the desired metadata property was available. In the case of OA, we counted the item if the OA status was not marked as closed. We have normalised the journal title to lowercase to aggregate the items.
3 Results and Discussion

### v3

@@ Abstract
OpenAlex is a promising open source of scholarly metadata, and competitor to established proprietary sources, such as the Web of Science and Scopus. As OpenAlex provides its data freely and openly, it permits researchers to perform bibliometric studies that can be reproduced in the community without licensing barriers. However, as OpenAlex is a rapidly evolving source and the data contained within is expanding and also quickly changing, the question naturally arises as to the trustworthiness of its data. In this report, we will study the reference coverage and selected metadata within each database and compare them with each other to help address this open question in bibliometrics. In our large-scale study, we demonstrate that, when restricted to a cleaned dataset of 16.8 million recent publications shared by all three databases, OpenAlex has average source reference numbers and internal coverage rates comparable to both Web of Science and Scopus. We further analyse the metadata in OpenAlex, the Web of Science and Scopus by journal, finding a similarity in the distribution of source reference counts in the Web of Science and Scopus as compared to OpenAlex. We also demonstrate that the comparison of other core metadata covered by OpenAlex shows mixed results when broken down by journal, capturing more ORCID identifiers, fewer abstracts and a similar number of Open Access status indicators per article when compared to both the Web of Science and Scopus.
@@
Bibliometrics, Open Scholarly Metadata, Citation Analysis, Reference Coverage, Scholarly Databases, OpenAlex
@@ 1 Introduction
OpenAlex is a promising alternative to proprietary bibliometric data sources as its permissible licensing creates the potential to support a transformation of research practice towards reproducible bibliometrics. This is being realised in open research policies in academia, for example in December 2023, Sorbonne University has switched from using the Web of Science (WoS) and Clarivate bibliometric tools to OpenAlex and open-source tools. [^1] https://www.sorbonne-universite.fr/en/news/sorbonne-university-unsubscribes-web-science Reproducible bibliometric research is hardly possible with proprietary bibliometric data sources as their licensing terms rule out dissemination of data.
@@
As a widely used open source repository of scholarly metadata, OpenAlex has previously been the subject of research as to its suitability for a variety of bibliometric analyses, such as the review by [25], which comparatively analysed various APIs to bibliometric corpora, including API interoperability, characteristics and their use in research practice, and [2] who released a working paper on the migration of scholars which included a comparative study between Scopus and OpenAlex, limited to the coverage of scholars in Western and non-Western countries.
Although country affiliation metadata accuracy and completeness were found lacking in earlier versions [30], OpenAlex was recently considered suitable for countrywide analyses by [1]. Investigating diamond open access journals indexed in OpenAlex in comparision with both WoS and Scopus, [21] highlighted that OpenAlex journal indexing is more inclusive than that of WoS and Scopus. However, at this early stage of its development, OpenAlex is a highly dynamic data source whose characteristics change with each release. This paper also follows previous quantitative comparisons of citation coverage of traditional bibliometric databases. One such study examines a dataset of 2.5 thousand documents published in 2006 and approximately 3.1 million citations of these documents, found in the MAG, Google Scholar, WoS, Scopus and OpenCitations’ COCI databases, [15]. This study demonstrates relative coverage gaps in some subject areas in the MAG as compared to WoS and Scopus, and also overall that Google Scholar has the largest citation coverage as compared to the other databases.
This concurs with an earlier study also by [14], which also compared these databases to Google Scholar, where approximately 2.45 million citations from 2300 documents covering 252 subjects are compared for similarities in the citation coverage, similarly finding the citations in Google Scholar comprises a superset of those in WoS and Scopus.
Previous work has studied and compared bibliographic databases to better understand the limitations of different data sources. Comparing the MAG, Scopus, WoS and other databases, [29], argue for combining databases to allow for comprehensive coverage, taking into account the strengths and weaknesses of the different data sources. Similarly, other studies have focused on a journal coverage analysis of WoS, Scopus and Dimensions, such as the study of [24]. Furthermore tools such as Färber’s tool for comparing author records between databases [7], have been created and shared by the academic community to provide insights into the suitability and weaknesses of bibliometric databases for accurate bibliometrics in their current state.
@@
Therefore to enable bibliometricians to better understand the potentials and current limitations of OpenAlex, we compare OpenAlex with two major proprietary bibliometric data sources, WoS and Scopus. With our study, we wish to contribute to the question to what extent OpenAlex can serve as an adequate, (or even better) free alternative to established, proprietary databases for bibliometric research and reporting. Our specific research questions in this report are whether reference coverage of items differs between the three data sources, investigating this for the complete databases as well as for a sub-corpus of items present in all three databases, and whether and to what extent the coverage of some additional metadata fields, specifically abstracts, Open Researcher and Contributor IDs (ORCIDs), and Open Access status of items differ in all three data sources.
We are aware that these initial assessments are likely to change with further developments, as of writing twelve new snapshots of OpenAlex have been released - with new data added or modified regularly, so this report should be understood as reflecting the state as of late 2023. Since then at least 151 million new references have been added which was an increase of 7.61% while at least 750,000 records were deleted and over 3.4 million records were added. [^2] https://github.com/ourresearch/openalex-guts/blob/main/files-for-datadumps/standard-format/RELEASE_NOTES.txt
@@ 1.1 Reference Coverage
References are of central importance for bibliometric databases, as matching them to their target items forms the basis for the calculation of citation metrics. As a first step, we compare average reference counts between the three databases, whereby the basis of the comparison are the complete databases, then subsets of publications with the document type ‘article’, and a shared sub-corpus of publications covered by all three databases. Citation reference data can also be used for an indirect assessment of the coverage, i.e. the proportion of relevant research publications that are included in the database and accessible to users for analysis [24]. An insufficient or biased coverage of the relevant literature should rule out the use of a database for a particular study.
@@
There are different ways to determine the coverage of a database, for example, the comparison with external lists of relevant sources or publication lists of a sample of representative researchers of the studied fields. However, there is no general gold standard corpus and all external sources therefore bring their own biases and limitations. A relatively simple (and easily replicable and repeatable) way to study literature coverage is calculating the internal reference coverage of a database as a whole or in relation to grouping characteristics, such as disciplines, the literature of particular countries or language communities.
@@
The internal coverage is the proportion of those cited references of a publication set which are themselves covered as source items in the database, out of all cited references in the set. We refer to these as source references and in contrast, to references to items that are not themselves indexed in the database as non-source references (or references to non-source items).
@@
A more comprehensive introduction to this concept and an analysis of the internal coverage is available in Moed [16, Chapter 7] and van Raan [28]. The great advantage of this type of analysis is that one does not need any external data which may be difficult and costly to collect. This reliance on only the assessed data source itself is also the major disadvantage, as one is limited to the reference data as present in the assessed data source with all its contingencies. Therefore one cannot simply extrapolate from the coverage of cited literature to the coverage of literature segments that were never cited in the source data, possibly as a direct consequence of the source database’s selection criteria. These considerations show why internal reference coverage provides merely a partial and possibly source-biased measurement of coverage.
@@
Nevertheless, when comparing citation index databases, the differences in internal reference collection can be a useful guide. For example, one question that arises due to the much larger dimension of OpenAlex compared to WoS and Scopus is whether it thereby also has a higher internal coverage, i.e. a higher proportion of publications that are referenced and also indexed in the database compared to the other databases. There are no established guidelines for numerical values of coverage proportions required to allow reliable studies to be carried out. But for example, [16] analysed the combined ISI Citation Indexes (the predecessor of today’s Web of Science) and found that the coverage rate, which is the proportion of references from the 2002 source year that refer to ISI source journals, was highest for Molecular Biology and Biochemistry, at around 90%, followed by human-focused Biological Sciences, Chemistry, Clinical Medicine and Physics and Astronomy. It was vastly lower in the Arts and Humanities and intermediary in the Social Sciences, Mathematics and Engineering.
@@
When using this indicator to compare OpenAlex, WoS and Scopus, we are thus less interested in an evaluation in absolute values, but rather in assessing how OpenAlex performs in comparison to the two established bibliometric databases. Internal reference coverage depends on the size and possibly the disciplinary profile of a database as well as the accuracy of its reference matching procedure. As OpenAlex is actually much larger than Scopus and WoS (see Table 1) it could be expected that its internal reference coverage is at least not lower than those of the latter databases.
@@ 1.2 Open Metadata
The increasing discussion surrounding the open availability and quality of various types of scholarly metadata in bibliometrics is not limited to reference coverage, but expands to other metadata [26, 5, 4, 30]. For instance, the Initiative for Open Abstracts (I4OA) [^4] https://i4oa.org/ advocates open abstracts of scholarly works and calls on scholarly publishers to submit them to Crossref, [^3] https://www.crossref.org/ a Digital Object Identifier (DOI) registration agency. Similarly, scholarly publishers can use Crossref to share the funding information associated with the articles they publish. However, coverage analyses of Crossref suggest that not all publishers provide open scholarly metadata to Crossref [17, 12]. Another example of essential metadata is the use of ORCIDs to persistently identify authors, helping bibliometricians not only to disambiguate author names, but also to interlink different data from different sources based on the ORCID [9].
As open data sources are essential for OpenAlex, we will expand our analysis to compare abstracts, funding information and ORCID coverage at the journal level. Moreover, we will assess the coverage of open access status information between OpenAlex and the proprietary databases WoS and Scopus. In contrast to abstract and author information, all three databases use the same source, the open access discovery service Unpaywall, to retrieve open access status information [6].
2 Data and Methodology
In this section, we describe the data used in this study, and the reasoning for our choices of restrictions and subsets of this data. To enable a fair comparison between OpenAlex, and WoS and Scopus, we have created a ‘Shared Corpus’ containing records common to all three datasets based on an exact DOI match, which have been published between 2015 and 2022, where the DOI is unique to the record in all three databases, i.e. there are no multiple records with the same DOI. In the course of selecting records from the databases, it is ensured that publications only ever have one DOI assigned to the record. In a further step, the references of the publications in the Shared Corpus are restricted to those published between 1996 to 2022.
@@ 2 Data and Methodology
Table 1: Sizes of databases and of the Shared Corpus dataset, with the number of references contained in each dataset
The versions of the WoS, Scopus and OpenAlex databases used in this study are as follows. The WoS and Scopus data are snapshots from five indexes of the WoS Core Collection (Science Citation Index Expanded, Social Sciences Citation Index, Arts & Humanities Citation Index, Conference Proceedings Citation Index – Science and Conference Proceedings Citation Index – Social Sciences) starting from publication year 1980 and the Scopus database, both captured in April 2023. The OpenAlex database is the version released in August 2023, due in both cases to the versioning policy of our data host at FIZ Karlsruhe, the Leibniz Institute for Information Infrastructure.555www.fiz-karlsruhe.de
@@
As the Scopus database mainly contains items from 1996 onwards (although since 2015, pre-1996 cited references and backfiles of major publishers have been added, [^6] https://blog.scopus.com/posts/breaking-the-1996-barrier-scopus-adds-nearly-4-million-pre-1996-articles-and-more-than-83) and WoS and OpenAlex have had no such restriction, to avoid bias in the computation of source reference counts and internal coverage we further restrict references to those items published between 1996 and 2022.
We include a section on articles published 2015-2022 in Table 1 for all three databases, to illustrate the influence of the time restriction to the size of the Shared Corpus, and to give context to the DOI matching and deduplication work described in Section 2.1.
In the Scopus and WoS databases, pre-computed total ’reference counts’, are delivered by the data providers Elsevier and Clarivate, whereas ’source reference counts’ are calculated for each record by our data provider FIZ Karlsruhe. Both databases are expected to contain all references of a given publication, regardless of whether they refer to items contained within or not contained within their databases, i.e. whether they are source and non-source references respectively, and without a fixed time restriction.
In contrast, presently in OpenAlex there only exist source references (see the OpenAlex documentation [20] [^7] https://docs.openalex.org/api-entities/works/work-object#referenced_works – apart from a smaller segment of references to supposedly deleted items) and a ‘source reference count’ has been calculated by FIZ Karlsruhe in our database. This fact explains the empty values for the average total reference counts in Table 3. We therefore have to relate this number to the source reference counts taken from WoS and Scopus.
@@
To address the potential bias arising from the number of documents in the databases, we calculated the five-number summaries (median, standard deviation, maximum, minimum and inter-quartile range) in addition to the mean values for the number references per article. Despite the considerable variation in the number of references per article and number of documents in the databases, we did not observe substantial disparities in the distribution across the different data sources examined. Consequently, we have chosen to present the mean value.
Figure 1: Venn diagram of the intersection sizes of unique DOIs based in each database on exact DOI match (without deduplication, i.e. cases of DOIs that have been assigned to multiple papers are now kept in the sets), for records published between 2015 and 2022.
@@
To evaluate the reference and source reference coverage of WoS and Scopus against OpenAlex, we first used the reported reference counts and pre-calculated source reference counts as described in Section 2. The average total reference count and source reference count was computed for: each database, for records marked as ‘article’ (or comprising the document type ‘article’ alongside other type markings, in the case of Scopus and WoS) and for the Shared Corpus resulting from the DOI match (publication years 2015-2022).
These numbers were then checked by an independent calculation where the total number of references and records in each database were counted and the ratio was computed (‘references per record’), as reported in Table 4. Then for the final results, queries were created to calculate and average the number of references with reference publication year 1996 to 2022, and the number of references that are linked to source items and publication years 1996 to 2022.
@@ 2.1 DOI Match and Deduplication
When constructing the Shared Corpus as described in Section 2, we relied on the DOI as a unique identifier that we could use to combine the databases. This approach has its limitations, as explored in [27] which highlights the distribution of non-existent or duplicate DOIs in each corpus may vary by subject in the WoS and Scopus databases.
We excluded records without a DOI and records where more than one publication item is attributed to the same DOI – as we are virtually not able to decide which item is the correct one for a given DOI in the latter case. These duplicate records account for the removal of 39,481 publications (counted as distinct DOI) in addition to those resulting from the restriction to 2015-2022. This accounts for the difference between the size of the Shared Corpus and the nominal intersection of the three databases between 2015 and 2022.
@@ 2.1.1 Error Margins of the DOI Match
Records with a duplicate DOI or without a DOI were excluded from the DOI matching step in the construction of the Shared Corpus. In Table 2, which focuses on all publications in the three databases which are published between 2015 and 2022, it can be seen that Scopus has a significantly larger number of DOIs with multiple records associated with it. Altogether, OpenAlex has the greatest number of records without DOI, then WoS and Scopus.
As records without a DOI are not matched in our analyses there is a significant underestimation of the total size of the databases as portrayed in Figure 1, similarly records which have a shared DOIs are counted once.
@@
2.2 Metadata Coverage
To determine metadata coverage (as detailed in Section 3.3), we also used the Shared Corpus as described at the beginning of Section 2. Here, we restrict to publication items published in journals. For this purpose, the publication type categorisations of Web of Science and Scopus were used and the OpenAlex publications were assigned to these via the DOI comparison of the Shared Corpus, so that OpenAlex could be compared bilaterally with the other two databases. We then specifically compared the coverage of abstracts, funding information, ORCIDs and Open Access (OA) status information by assessing whether items have (at least one) of these and aggregated by journal, that is, for each journal, a publication record was counted if the desired metadata property was available. In the case of OA, we counted the item if the OA status was not marked as closed. We have normalised the journal title to lowercase to aggregate the items.
3 Results